OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y=-1.7x+8.5

OpenStudy (anonymous):

A. 17x+10y=85 B. -17x+10y=85 C. -17x+10y=-85 D. 17x+10y=-85

hartnn (hartnn):

example : if a = 0.4 10 a = 4 If you want to shift the decimal point to right, multiply by 10

OpenStudy (texaschic101):

first multiply everything in the equation by 10 to get rid of the decimal. Try that first...don't forget to multiply EVERYTHING by 10. So what did you end up with ?

OpenStudy (anonymous):

\[10y=-17x+85\]

OpenStudy (texaschic101):

Let me set it up and maybe it will be easier... (10) y = (10)-1.7x + (10)8.5.......do the math and what do you end up with ?

OpenStudy (texaschic101):

correct... 10y = -17x + 85...but we need it in standard form Ax + By = C so add 17x to both sides and what do you get ?

OpenStudy (anonymous):

10y+17x=85

OpenStudy (anonymous):

well 17x+10y=85

OpenStudy (texaschic101):

yep...or 17x + 10y = 85...answer choice 1

OpenStudy (anonymous):

yay thank you so much!!

OpenStudy (texaschic101):

sure thing :)